### v1

@@ 3.1. Impulsive noise
In general, a noise vector or function $\xi:\mathbb{M}_{h}\to\mathbb{R}$ is called impulsive, if $\lvert\xi\rvert$ is large on a small part of $\mathbb{M}_{h}$, and small or even zero elsewhere. This type of noise occurs e.g. in measurements with malfunctioning receivers, digital image acquisition with faulty memory locations, or in powerline communication systems, see also [CHN04, CJ12].
@@
Let $\left(\mathbb{M}_{h},\Sigma,\nu\right)$ be a (possibly discrete) measure space on which the observations live. A noise function $\xi\in\mathbf{L}^{1}\left(\mathbb{M}_{h}\right)$ is called impulsive with parameters $\varepsilon,\eta\geq 0$, if there exists $\mathbb{M}_{\mathrm{c}}\in\Sigma$ such that
@@
with an impulsive noise function $\xi=\xi\left(\varepsilon,\eta\right)\in\mathbf{L}^{1}\left(\mathbb{M}_{h}\right)$. Note, that we do not require any (direct) relation between $g^{\mathrm{obs}}$ and $g^{\dagger}=F\left(u^{\dagger}\right)$ here, but only between $g^{\mathrm{obs}}$ and the natural discretization $F_{h}\left(u^{\dagger}\right)$. The misfit between $F_{h}\left(u^{\dagger}\right)$ and $g^{\dagger}$ will be bounded below by assuming properties of $F_{h}$ as an approximation of $F$.
@@ 3.2. Resulting error bounds
Suppose the noise model (3.1) with an impulsive noise $\xi=\xi\left(\varepsilon,\eta\right)\in\mathbf{L}^{1}\left(\mathbb{M}_{h}\right)$, and let $F_{h}$ satisfy (3.2). Then the overall data model is $\left(\mathbf{L}^{\infty}\left(\mathbb{M}\right),\mathcal{Y}_{\mathrm{d}}\right)$-admissible with $\eta_{0}=2\epsilon$, $\eta_{1}=2C_{\omega}\eta$ and $\eta_{2}=C_{\omega}h^{\omega}$.
@@
With this estimate, we are now in position to specify the result from Theorem 2.4, cf. also [HW22, Cor. 6.1]
@@
Suppose the noise model (3.1) with an impulsive noise $\xi=\xi\left(\varepsilon,\eta\right)\in\mathbf{L}^{1}\left(\mathbb{M}_{h}\right)$, and let $F_{h}$ satisfy (3.2). Let furthermore the variational source condition (2.2) hold true, $F:D(F)\to\mathbf{W}^{k,p}\left(\mathbb{M}\right)$ be Lipschitz continuous with $k>\frac{d}{p}$ as in (2.7) and suppose $F$ is $\left(\mathcal{Y}_{d},\gamma,\ell\right)$-smoothing. Then there exists a constant $C$ independent of $\alpha,\varepsilon,\eta$ and $h$ such that any minimizer of (1.2) (if existent) satisfies the error bound
@@
Note, that the above result can be generalized to exponentially smoothing operators, i.e. to the case that $F:D(F)\to\mathbf{W}^{k,p}\left(\mathbb{M}\right)$ is Lipschitz for any choice of $k\in\mathbb{N}$, see [KWH16]. There, similar bounds with $\eta^{\vartheta+1}$ replaced by $\tilde{\gamma}\left(\eta\right)$ have been derived, where the function $\tilde{\gamma}$ tends faster to zero than any polynomial. For the backwards heat equation, one e.g. obtains $\tilde{\gamma}\left(\delta\right)\sim\exp\left(-\frac{1}{\delta^{2}}\right)$, and for three-dimensional satellite gradiometry $\tilde{\gamma}\left(\delta\right)\sim\delta^{-\frac{5}{2}}R^{-\sqrt{\frac{4\pi}{\delta}-4}}$, where $R>1$ is the radius of the measurement shell. However, for simplicity, we restrict ourselves to the finitely smoothing case as discussed above in this work.
@@ 3.3. Results for finitely smoothing operators under classical smoothness conditions
Let us now derive error bounds in the setting of Example 2.1 for generalized Tikhonov regularization with $\mathbf{L}^{1}$ data fidelity term. Note, that Corollary 3.3 is not directly applicable, since Example 2.1 only implies the variational source condition (2.4) with $\mathbf{L}^{2}$ data fidelity term. We therefore exploit ideas from [KWH16, Prop. 2.2] and mimic the proofs from [HW22] to obtain the following:
@@
    $F:D(F)\subset\mathbf{L}^{2}\left(\Omega\right)\to\mathbf{W}^{k,p}$ is Lipschitz continuous w.r.t. $\ell$ where $k>\frac{d}{p}$, this is
@@
Then, provided $\alpha,\varepsilon,\eta$ and $h$ are sufficiently small, any minimizer $\hat{u}_{\alpha}$ of (1.2) (if existent) obeys the error bound
@@
In the proof, we will make use of Young’s inequality with weights, this is
@@
which holds true for all $x,y\geq 0$, $w>0$, and $q\geq 1$ with $\frac{1}{q}+\frac{1}{q^{\prime}}=1$, see [HW22, Eq. (3.1)].
@@
for any $\delta\in(0,\delta_{0})$. Optimizing this bound in $\delta$ (at least in the limit $\alpha\searrow 0$ where the corresponding $\delta$ will satisfy $\delta<\delta_{0}$) yields (3.6). ∎
@@
    If $F$ is at least $a$-smoothing for some $a\in\mathbb{N}$, $a>\frac{d}{2}$, i.e. (2.3) is complemented by (2.8), then Assumption (A4) is automatically satisfied as well with $k=a$, $p=2$.
@@
    Compared to the approach from [KWH16], where the $\mathrm{res}_{\infty}$-term in $\mathrm{res}_{2}^{2}\leq\mathrm{res}_{\infty}\mathrm{res}_{1}$ is bounded using $\mathrm{res}_{\infty}\leq C$ uniformly, we get a different exponent of $\alpha$ in (3.6). By the uniform bound one would obtain $\exp_{\mathrm{unif}}:=\frac{s}{a}$, whereas above we derived $\exp_{\mathrm{fine}}:=\frac{2s(2\theta+1)}{(2\vartheta+1)(a-s)+a}$. In the situation of (b), i.e. if $F$ is at least $a$-smoothing in the sense of (2.8) and at most $a$-smoothing in the sense of (2.3), we can choose $p=2$ and $k=a$, which results in $2\vartheta+1=\frac{2a}{d}$, so that $\exp_{\mathrm{fine}}=\frac{4s}{2a-2s+d}>\exp_{\mathrm{unif}}$, i.e. our new bound is an improvement.
@@ 4.1. General setup
where $\xi_{i}$ for $i=1,\ldots,n$ are i.i.d. centered random variables and $x_{i}\in\mathbb{M}$ are sampling points for $i=1,\ldots,n$. Here we assume w.l.o.g. that $g^{\dagger}$ can be interpreted as a continuous function, which is e.g. the case under Assumption (A4) from Theorem 3.5. To include this setup into our previous model, we set
@@
Note, that (3.2b) is trivially fulfilled in this setup, and (3.2a) holds true e.g. with $\mathcal{Y}_{\mathrm{d}}=\mathbf{L}^{\infty}\left(\mathbb{M}\right)$ as soon as $R(F)$ consists of sufficiently smooth functions, as it is typically the case in inverse problems.
Let us now take a threshold $\lambda>0$. If we define
@@
(4.2) $\varepsilon=\left\|\xi\right\|_{\mathbf{L}^{1}\left(\mathbb{M}_{h}\setminus\mathbb{M}_{\mathrm{c}}\right)}=\frac{1}{n}\sum_{k\notin\mathbb{M}_{\mathrm{c}}}|\xi_{k}|,\qquad\eta=\nu\left(\mathbb{M}_{\mathrm{c}}\right)=\frac{\#\mathbb{M}_{\mathrm{c}}|\mathbb{M}|}{n}.$
Note, that $\eta$ and $\varepsilon$ are random variables as well.
@@
Let $\xi_{i},i=1,\ldots,n$ i.i.d. centered random variables with cdf $F_{\xi}$. For the quantities $\varepsilon$ and $\eta$ defined in (4.2), it holds that
$\mathbb{E}[\varepsilon]\leq\sqrt{F_{\xi}\left(\lambda\right)-F_{\xi}\left(-\lambda\right)}\sqrt{\mathbb{V}\left[\xi_{1}\right]}\qquad\text{and}\qquad\mathbb{E}[\eta^{r}]\leq|\mathbb{M}|^{r}\left(1-F_{\xi}\left(\lambda\right)+F_{\xi}\left(-\lambda\right)\right)$
@@
Recall $\mathbb{M}_{\mathrm{i}}:=\mathbb{M}_{h}\setminus\mathbb{M}_{\mathrm{c}}$. We compute using Cauchy-Schwarz that
$\mathbb{E}[\varepsilon]=\mathbb{E}\left[\frac{1}{n}\sum_{k\in\mathbb{M}_{\mathrm{i}}}|\xi_{k}|\right]=\frac{1}{n}\mathbb{E}\left[\sum_{i=1}^{n}1|_{\{|\xi_{i}|\leq\lambda\}}\left\|\xi_{i}\right\|\right]=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left[1|_{\{|\xi_{i}|\leq\lambda\}}\left|\xi_{i}\right|\right]\\ =\mathbb{E}\left[1|_{\{|\xi_{1}|\leq\lambda\}}\left|\xi_{1}\right|\right]\leq\sqrt{\mathbb{E}\left[1|_{\{|\xi_{1}|\leq\lambda\}}\right]}\sqrt{\mathbb{E}\left[\left|\xi_{1}\right|^{2}\right]}=\sqrt{\mathbb{P}\left[\left|\xi_{1}\right|\leq\lambda\right]}\sqrt{\mathbb{V}\left[\xi_{1}\right]}\\ =\sqrt{F_{\xi}\left(\lambda\right)-F_{\xi}\left(-\lambda\right)}\sqrt{\mathbb{V}\left[\xi_{1}\right]}$
@@
Note, that not only the moments of $\varepsilon$ and $\eta$ can be bounded, but it is also possible to derive concentration results by standard techniques. Therefore, all the following considerations would also allow for concentration results.
@@ 4.2. Gaussian case and optimal rates of convergence
Let $\xi_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}\left(0,\sigma^{2}\right)$ for $i=1,\ldots,n$ with $\sigma>0$. Then, for any $r\geq 1$, the moments of $\varepsilon$ and $\eta$ as defined in (4.2) can be bounded by
$\mathbb{E}[\varepsilon]\leq\sigma\sqrt{1-2\Phi\left(-\frac{\lambda}{\sigma}\right)},\qquad\mathbb{E}[\eta^{r}]\leq 2|\mathbb{M}|^{r}\Phi\left(-\frac{\lambda}{\sigma}\right)$
@@
In the setting of Theorem 3.5, grant Assumptions (A2)–(A6) and let $g^{\mathrm{obs}}$ be given as in (4.1) with $\xi_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}\left(0,\sigma^{2}\right)$ for $i=1,\ldots,n$ with $\sigma>0$. Then, for sufficiently small $\alpha,\sigma$ and $h$, any minimizer $\hat{u}_{\alpha}$ of (1.2) (if existent) obeys the bound
@@
Let us now derive rates of convergence from this result. Therefore, we will assume that the discretization error, encoded in the term $h^{2\omega}\left(\gamma\left(3h^{\omega}\right)\right)^{2}$ is negligible compared to the statistical noise in the observations. Note, that this is the typical situation in Inverse Problems, where $F$ is smoothing and hence both the exponent $\omega$ in (3.2) will be comparably large and the function $\gamma$ will be decaying fast. In this situation, we only have to treat the bound
@@
In the setting of Theorem 3.5, grant Assumptions (A2)–(A6) and let $g^{\mathrm{obs}}$ be given as in (4.1) with $\xi_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}\left(0,\sigma^{2}\right)$ for $i=1,\ldots,n$ with $\sigma>0$. Suppose that the discretization error $h^{2\omega}\left(\gamma\left(3h^{\omega}\right)\right)^{2}$ is negligible. Then, for an optimally chosen $\alpha=\alpha\left(\sigma,\lambda\right)$, any minimizer $\hat{u}_{\alpha}$ of (1.2) (if existent) satisfies
@@
Note, that this shows a principle bound
@@
In the setting of Theorem 3.5, grant Assumptions (A2)–(A6) and let $g^{\mathrm{obs}}$ be given as in (4.1) with $\xi_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}\left(0,\sigma^{2}\right)$ for $i=1,\ldots,n$ with $\sigma>0$. Suppose that the discretization error $h^{2\omega}\left(\gamma\left(3h^{\omega}\right)\right)^{2}$ is negligible. Then, for an optimally chosen $\alpha$, any minimizer $\hat{u}_{\alpha}$ of (1.2) (if existent) satisfies
@@
We aim to choose $\lambda>0$ such that both terms in the maximum in (4.3) equal each other. Since we want the right-hand side to go to 0 as $\sigma\searrow 0$, this requires $z=\Phi\left(-\frac{\lambda}{\sigma}\right)\to 0$ as well. Consequently, the $\sqrt{1-2z}$-term in the first argument of the maximum can just be ignored, as it will be bounded for $z\to 0$. Since, for every $\sigma>0$, there exists a $\lambda>0$ (with the asymptotics $\lambda\to\infty$ as $\sigma\searrow 0$) with $\sigma^{\frac{\theta}{\theta+1}}=z=\Phi\left(-\frac{\lambda}{\sigma}\right)$, this shows that both terms in the maximum (4.3) can be equaled (asymptotically), yielding the claimed bound. ∎
@@
In the setting of Theorem 3.5, grant Assumptions (A3), (A5) and (A6) and assume that $F$ is $a$-times smoothing for some $a\in\mathbb{N}$, $a>d/2$ as in (2.3) and (2.8). Let $g^{\mathrm{obs}}$ be given as in (4.1) with $\xi_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}\left(0,\sigma^{2}\right)$ for $i=1,\ldots,n$ with $\sigma>0$. Suppose that the discretization error $h^{2\omega}\left(\gamma\left(3h^{\omega}\right)\right)^{2}$ is negligible. Then, for an optimally chosen $\alpha$, any minimizer $\hat{u}_{\alpha}$ of (1.2) (if existent) satisfies

### v2

@@ 3.1. Impulsive noise
In general, a noise vector or function $\xi:\mathbb{M}_{h}\to\mathbb{R}$ is called impulsive if $\lvert\xi\rvert$ is large on a small part of $\mathbb{M}_{h}$, and small or even zero elsewhere. This type of noise occurs e.g. in measurements with malfunctioning receivers, digital image acquisition with faulty memory locations, or in powerline communication systems, see also [CHN04, CJ12].
@@
Let $\left(\mathbb{M}_{h},\Sigma,\nu\right)$ be a (possibly discrete) measure space on which the observations are defined. A noise function $\xi\in\mathbf{L}^{1}\left(\mathbb{M}_{h}\right)$ is called impulsive with parameters $\varepsilon,\eta\geq 0$, if there exists $\mathbb{M}_{\mathrm{c}}\in\Sigma$ such that
@@
with an impulsive noise function $\xi=\xi\left(\varepsilon,\eta\right)\in\mathbf{L}^{1}\left(\mathbb{M}_{h}\right)$. Note that we do not require any (direct) relation between $g^{\mathrm{obs}}$ and $g^{\dagger}=F\left(u^{\dagger}\right)$ here, but only between $g^{\mathrm{obs}}$ and the natural discretization $F_{h}\left(u^{\dagger}\right)$. The misfit between $F_{h}\left(u^{\dagger}\right)$ and $g^{\dagger}$ will be bounded in the subsection below by assuming properties of $F_{h}$ as an approximation of $F$.
@@ 3.2. Resulting error bounds
Suppose the noise model (3.1) with an impulsive noise $\xi=\xi\left(\varepsilon,\eta\right)\in\mathbf{L}^{1}\left(\mathbb{M}_{h}\right)$, and let $F_{h}$ satisfy (3.2). Then the overall data model is $\left(\mathbf{L}^{\infty}\left(\mathbb{M}\right),\mathcal{Y}_{\mathrm{d}}\right)$-admissible (i.e. $\mathcal{Y}_{\mathrm{d}}^{1}=\mathbf{L}^{\infty}\left(\mathbb{M}\right)$ and $\mathcal{Y}_{\mathrm{d}}^{2}=\mathcal{Y}_{\mathrm{d}}$) with $\eta_{0}=2\epsilon$, $\eta_{1}=2C_{\omega}\eta$ and $\eta_{2}=C_{\omega}h^{\omega}$.
@@
With this estimate, we are now in a position to specify the result from Theorem 2.4, cf. also [HW22, Cor. 6.1]
@@
Suppose the noise model (3.1) with an impulsive noise $\xi=\xi\left(\varepsilon,\eta\right)\in\mathbf{L}^{1}\left(\mathbb{M}_{h}\right)$, and let $F_{h}$ satisfy (3.2). Let furthermore the variational source condition (2.2) hold true, $F:D(F)\to\mathbf{W}^{k,p}\left(\mathbb{M}\right)$ be Lipschitz continuous with $k>\frac{d}{p}$ as in (2.7) and suppose $F$ is $\left(\mathcal{Y}_{d},\gamma,\ell\right)$-smoothing. Assume $r>1$. Then there exists a constant $C$ independent of $\alpha,\varepsilon,\eta$ and $h$ such that any minimizer of (1.2) (if it exists) satisfies the error bound
@@
Note that the above result can be generalized to exponentially smoothing operators, i.e. to the case that $F:D(F)\to\mathbf{W}^{k,p}\left(\mathbb{M}\right)$ is Lipschitz for any choice of $k\in\mathbb{N}$, see [KWH16]. There, similar bounds with $\eta^{\vartheta+1}$ replaced by $\tilde{\gamma}\left(\eta\right)$ have been derived, where the function $\tilde{\gamma}$ tends faster to zero than any polynomial. For the backwards heat equation, one e.g. obtains $\tilde{\gamma}\left(\delta\right)\sim\exp\left(-\frac{1}{\delta^{2}}\right)$, and for three-dimensional satellite gradiometry $\tilde{\gamma}\left(\delta\right)\sim\delta^{-\frac{5}{2}}R^{-\sqrt{\frac{4\pi}{\delta}-4}}$, where $R>1$ is the radius of the measurement shell. However, for simplicity, we restrict ourselves to the finitely smoothing case as discussed above in this work.
@@ 3.3. Results for finitely smoothing operators under classical smoothness conditions
Let us now derive error bounds in the setting of Example 2.1 for generalized Tikhonov regularization with $\mathbf{L}^{1}$ data fidelity term. Note that Corollary 3.3 is not directly applicable, since Example 2.1 only implies the variational source condition (2.4) with $\mathbf{L}^{2}$ data fidelity term. We therefore exploit ideas from [KWH16, Prop. 2.2] and mimic the proofs from [HW22] to obtain the following:
@@
    $F:D(F)\subset\mathbf{L}^{2}\left(\Omega\right)\to\mathbf{W}^{k,p}\left(\mathbb{M}\right)$ is Lipschitz continuous w.r.t. $\ell$ where $k>\frac{d}{p}$, that is,
@@
Then, provided $\alpha,\varepsilon,\eta$ and $h$ are sufficiently small, any minimizer $\hat{u}_{\alpha}$ of (1.2) (if it exists) obeys the error bound
@@
In the proof, we will make use of Young’s inequality with weights, that is,
@@
which holds true for all $x,y\geq 0$, $w>0$, and $q>1$ with $\frac{1}{q}+\frac{1}{q^{\prime}}=1$, see [HW22, Eq. (3.1)].
@@
for any $\delta\in(0,\delta_{0})$. Optimizing this bound with respect to $\delta$ (at least in the limit $\alpha\searrow 0$ where the corresponding $\delta$ will satisfy $\delta<\delta_{0}$) yields (3.6). ∎
@@
    If $F$ is at least $a$-smoothing for some $a\in\mathbb{N}$, $a>\frac{d}{2}$, i.e. (2.3) is complemented by (2.8), then Assumption (A4) is automatically satisfied as well with $k=a$, $p=2$. As mentioned before, since $a>\frac{d}{2}$, the image of $F$ is contained in $\mathbf{L}^{\infty}\left(\mathbb{M}\right)$ and hence by boundedness of $\mathbb{M}$, the operator $F$ can be considered as a map into $\mathbf{L}^{1}\left(\mathbb{M}\right)$.
@@
    Compared to the approach from [KWH16], where the $\mathrm{res}_{\infty}$-term in $\mathrm{res}_{2}^{2}\leq\mathrm{res}_{\infty}\mathrm{res}_{1}$ is bounded using $\mathrm{res}_{\infty}\leq C$ uniformly, we get a different exponent of $\alpha$ in (3.6). By the uniform bound one would obtain $\exp_{\mathrm{unif}}:=\frac{s}{a}$, whereas above we derived $\exp_{\mathrm{fine}}:=\frac{2s(2\theta+1)}{(2\vartheta+1)(a-s)+a}$. In the situation of (b), i.e. if $F$ is at least $a$-smoothing in the sense of (2.8) and at most $a$-smoothing in the sense of (2.3), we can choose $p=2$ and $k=a$, which results in $2\vartheta+1=\frac{2a}{d}$, so that $\exp_{\mathrm{fine}}=\frac{4s}{2a-2s+d}>\exp_{\mathrm{unif}}$, and in view of the additional requirement $a>d/2$ (see (b)), our new bound is an improvement.
@@ 4.1. General setup
where $\xi_{i}$, $i=1,\ldots,n$, are i.i.d. centered random variables and $x_{i}\in\mathbb{M}$ are sampling points for $i=1,\ldots,n$. Here we assume w.l.o.g. that $g^{\dagger}$ can be interpreted as a continuous function, which is e.g. the case under Assumption (A4) from Theorem 3.5. To include this setup into our previous model, we set
@@
Note that (3.2b) is trivially fulfilled in this setup, and (3.2a) holds true e.g. with $\mathcal{Y}_{\mathrm{d}}=\mathbf{L}^{\infty}\left(\mathbb{M}\right)$ as soon as $R(F)$ consists of sufficiently smooth functions, as is typically the case in inverse problems.
Consider a threshold $\lambda>0$. If we define
@@
(4.2) $\varepsilon=\left\|\xi\right\|_{\mathbf{L}^{1}\left(\mathbb{M}_{h}\setminus\mathbb{M}_{\mathrm{c}}\right)}=\frac{|\mathbb{M}|}{n}\sum_{k\notin\mathbb{M}_{\mathrm{c}}}|\xi_{k}|,\qquad\eta=\nu\left(\mathbb{M}_{\mathrm{c}}\right)=\frac{\#\mathbb{M}_{\mathrm{c}}|\mathbb{M}|}{n}.$
Note that $\eta$ and $\varepsilon$ are random variables as well.
@@
Let $\xi_{i},i=1,\ldots,n$, i.i.d. centered random variables with cdf $F_{\xi}$. For the quantities $\varepsilon$ and $\eta$ defined in (4.2), it holds that
$\mathbb{E}[\varepsilon]\leq|\mathbb{M}|\sqrt{F_{\xi}\left(\lambda\right)-F_{\xi}\left(-\lambda\right)}\sqrt{\mathbb{V}\left[\xi_{1}\right]}\quad\text{and}\quad\mathbb{E}[\eta^{r}]\leq|\mathbb{M}|^{r}\left(1-F_{\xi}\left(\lambda\right)+F_{\xi}\left(-\lambda\right)\right)$
@@
Recall $\mathbb{M}_{\mathrm{i}}:=\mathbb{M}_{h}\setminus\mathbb{M}_{\mathrm{c}}$. We compute, using Cauchy-Schwarz, that
$\mathbb{E}[\varepsilon]=\mathbb{E}\left[\frac{|\mathbb{M}|}{n}\sum_{k\in\mathbb{M}_{\mathrm{i}}}|\xi_{k}|\right]=\frac{|\mathbb{M}|}{n}\mathbb{E}\left[\sum_{i=1}^{n}1|_{\{|\xi_{i}|\leq\lambda\}}\left\|\xi_{i}\right\|\right]=\frac{|\mathbb{M}|}{n}\sum_{i=1}^{n}\mathbb{E}\left[1|_{\{|\xi_{i}|\leq\lambda\}}\left|\xi_{i}\right|\right]\\ =|\mathbb{M}|\mathbb{E}\left[1|_{\{|\xi_{1}|\leq\lambda\}}\left|\xi_{1}\right|\right]\leq|\mathbb{M}|\sqrt{\mathbb{E}\left[1|_{\{|\xi_{1}|\leq\lambda\}}\right]}\sqrt{\mathbb{E}\left[\left|\xi_{1}\right|^{2}\right]}=|\mathbb{M}|\sqrt{\mathbb{P}\left[\left|\xi_{1}\right|\leq\lambda\right]}\sqrt{\mathbb{V}\left[\xi_{1}\right]}\\ =|\mathbb{M}|\sqrt{F_{\xi}\left(\lambda\right)-F_{\xi}\left(-\lambda\right)}\sqrt{\mathbb{V}\left[\xi_{1}\right]}$
@@
Note that not only the moments of $\varepsilon$ and $\eta$ can be bounded, but it is also possible to derive concentration results by standard techniques. Therefore, all the following considerations would also allow for concentration results.
@@ 4.2. Gaussian case and optimal rates of convergence
Let $\xi_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}\left(0,\sigma^{2}\right)$ for $i=1,\ldots,n$ with $\sigma^{2}>0$. Then, for any $r\geq 1$, the moments of $\varepsilon$ and $\eta$ as defined in (4.2) can be bounded by
$\mathbb{E}[\varepsilon]\leq|\mathbb{M}|\sigma\sqrt{1-2\Phi\left(-\frac{\lambda}{\sigma}\right)},\qquad\mathbb{E}[\eta^{r}]\leq 2|\mathbb{M}|^{r}\Phi\left(-\frac{\lambda}{\sigma}\right)$
@@
In the setting of Theorem 3.5, grant Assumptions (A2)–(A6) and let $g^{\mathrm{obs}}$ be given as in (4.1) with $\xi_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}\left(0,\sigma^{2}\right)$ for $i=1,\ldots,n$ with $\sigma>0$. Then, for sufficiently small $\alpha,\sigma$ and $h$, any minimizer $\hat{u}_{\alpha}$ of (1.2) (if it exists) obeys the bound
@@
Let us now derive rates of convergence from this result. Therefore, we will assume that the discretization error, encoded in the term $h^{2\omega}\left(\gamma\left(3h^{\omega}\right)\right)^{2}$ is negligible compared to the statistical noise in the observations. Note that this is the typical situation in Inverse Problems, where $F$ is smoothing and hence both the exponent $\omega$ in (3.2) will be comparatively large and the function $\gamma$ will decay fast. In this situation, we only have to treat the bound
@@
In the setting of Theorem 3.5, grant Assumptions (A2)–(A6) and let $g^{\mathrm{obs}}$ be given as in (4.1) with $\xi_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}\left(0,\sigma^{2}\right)$ for $i=1,\ldots,n$ with $\sigma>0$. Suppose that the discretization error $h^{2\omega}\left(\gamma\left(3h^{\omega}\right)\right)^{2}$ is negligible. Then, for an optimally chosen $\alpha=\alpha\left(\sigma,\lambda\right)$, any minimizer $\hat{u}_{\alpha}$ of (1.2) (if it exists) satisfies
@@
Note that this shows (in principle) a bound
@@
In the setting of Theorem 3.5, grant Assumptions (A2)–(A6) and let $g^{\mathrm{obs}}$ be given as in (4.1) with $\xi_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}\left(0,\sigma^{2}\right)$ for $i=1,\ldots,n$ with $\sigma>0$. Suppose that the discretization error $h^{2\omega}\left(\gamma\left(3h^{\omega}\right)\right)^{2}$ is negligible. Then, for an optimally chosen $\alpha$, any minimizer $\hat{u}_{\alpha}$ of (1.2) (if it exists) satisfies
@@
We aim to choose $\lambda>0$ such that both terms in the maximum in (4.3) equal each other. Since we want the right-hand side to go to 0 as $\sigma\searrow 0$, this requires $z=\Phi\left(-\frac{\lambda}{\sigma}\right)\to 0$ as well. Consequently, the $\sqrt{1-2z}$-term in the first argument of the maximum can just be ignored, as it will be bounded for $z\to 0$. Since, for every $\sigma>0$, there exists a $\lambda>0$ (with the asymptotics $\lambda\to\infty$ as $\sigma\searrow 0$) with $\sigma^{\frac{\theta}{\theta+1}}=z=\Phi\left(-\frac{\lambda}{\sigma}\right)$, this shows that both terms in the maximum (4.3) can be equated (asymptotically), yielding the claimed bound. ∎
@@
In the setting of Theorem 3.5, grant Assumptions (A3), (A5) and (A6) and assume that $F$ is $a$-times smoothing for some $a\in\mathbb{N}$, $a>d/2$ as in (2.3) and (2.8). Let $g^{\mathrm{obs}}$ be given as in (4.1) with $\xi_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}\left(0,\sigma^{2}\right)$ for $i=1,\ldots,n$ with $\sigma>0$. Suppose that the discretization error $h^{2\omega}\left(\gamma\left(3h^{\omega}\right)\right)^{2}$ is negligible. Then, for an optimally chosen $\alpha$, any minimizer $\hat{u}_{\alpha}$ of (1.2) (if it exists) satisfies